OpenStudy (anonymous):

There are seven ice skaters in a competition. How many different possibilities are there for the top three places? (A different ordering of the same three skaters counts as a different possibility.

OpenStudy (anonymous):

\[7*3=21\]

OpenStudy (kropot72):

There are 7 possibilities for first place. For each of the 7 possibilities for first place there are 6 possibilities for second place and for each of the 7 * 6 possibilities for first and second place there are 5 possibilities for third place. Therefore the total number of possibilities for the top three places is found from \[7\times 6\times 5\]